OpenStudy (anonymous):

vertex: (0,0) directrix: y=-1 a. x=4y^2 b. x=1/4y^2 c.y=4x^2 d.y=1/4x^2

OpenStudy (kohai):

Are you trying to find the equation for this?

OpenStudy (anonymous):

yes. Thank you. haha

OpenStudy (kohai):

The vertex form of a parabola opening up or down, y=a(x-h)^2+k where (h,k) is the vertex The standard form is (x-h)^2+4p(y-k) where the focus is (h,k + p) y = (1/4)x^2, Vertex(0,0), focus(0,1) Directrix y=-1

OpenStudy (anonymous):

The directrix is as y=-1 so it is opening up, y=(1/4p)x^2 p= distance from vertex to focus/directrix since vertex is at the origin p=1 so y=(1/4)x^2

OpenStudy (anonymous):

thank you both! :)

OpenStudy (kohai):

You're welcome :)